# PsycoLLM: Enhancing LLM for Psychological Understanding and Evaluation

### v1

@@ 3.1 Dataset Collection and Preprocessing
In order to accurately evaluate the model performance in the psychology domain, we introduce a benchmark that psychological professional individuals need to master. The data is sourced from publicly available examination questions. We follow a standardized data preprocessing procedure. For some data, we first need to perform optical character recognition (OCR) to convert images to text. After that, we invite several students to manually review the collected data to ensure its quality and consistency with the original document. This process involves rectifying formatting errors, eliminating duplicate questions, and rectifying any instances of garbled characters. Our proposed psychological benchmark draws inspiration from the format of the most authoritative psychological counseling examination in China, comprising two primary components: objective questions and subjective questions. Note that the objective questions mainly comprise two types of multiple-choice questions (MCQs): MCQ with only a single correct option (SMCQ) and MCQ with multiple correct options (MMCQ). The content of benchmark can be categorized into three sections: professional ethics, theoretical proficiency, and case analysis. The professional ethics section generally encompasses general principles such as honesty, integrity, professional responsibility, and social responsibility. Theoretical proficiency typically pertains to fundamental theories and concepts in psychology, including cognitive psychology, developmental psychology, and related fields. Both the professional ethics and theoretical proficiency sections contain only MCQs, designed to evaluate the mastery of knowledge within the psychological domain. For case analysis, it includes both MCQs and iterative questioning based on given case, aiming to better evaluate performance in real-world scenarios. The former focuses on assessing the grasp of professional knowledge related to the case, while the latter offers a more open-ended assessment of the model’s capability in practical utility. Finally, we obtained a total of 3863 MCQs and 20 cases with 100 iterative questions. The statistics of the benchmark are shown in Table 2, detailing the distribution of questions across various categories. The questions are classified into three main categories: Level 2 Psychological Counselor Exam, Level 3 Psychological Counselor Exam, and Others. Note that, according to the official definitions, the Level 2 exam is considered more challenging than the Level 3 exam.
3.2 Evaluation Metric
@@
There are three elements in each instance of the case analysis. is the description of the given case, including the background, environment, etc. of the seeker, is the question according to the case, such as the diagnosis, reason and treatment for the case and is the reference answer for each questions. For QA in case analysis, we use the models to generate the response according to the given and . We evaluate the response from model with the corresponding references in terms of automatic evaluation. We follow existing text generation metric, including ROUGE-1, ROUGE-L Lin (2004), BLEU-4 Papineni et al. (2002), BERTScore Zhang et al. (2019).
For SMCQ, we use the standard accuracy to evaluate the model performance. For MMCQ, we employ two types of accuracy metrics: standard accuracy and elastic accuracy. In elastic accuracy, if the predicted answer contains correct options without any incorrect ones but is incomplete, the elastic accuracy for a MMCQ is computed as follows:
where || indicates the count of choices in the answer.
4 Experimental Setting
4.1 Baseline
In our experiments, our experiment is implemented based on transformers222https://github.com/huggingface/transformers and LLaMA-Factory333https://github.com/hiyouga/LLaMA-Factory.
BaiChuan2 Yang et al. (2023): This is a generation of open-source large language models launched by Baichuan Intelligent Technology, which is trained on a high-quality corpus with 2.6 trillion tokens.
ChatGLM3 Du et al. (2022): This is a generation of pre-trained dialogue models jointly released by Zhipu AI and Tsinghua KEG.
LLaMA3-Chinese: This model uses LLama3 as the base model and is fine-tuned with Chinese instruction datasets.
Yi444https://github.com/01-ai/Yi: This open-source large language models is trained from scratch by developers at 01.AI.
Qwen (e.g., Qwen1.5-14B-Chat, Qwen-turbo): these are decoder-only transformer models with SwiGLU activation, RoPE, multi-head attention, developed by Alibaba Cloud.
GPT-3.5-turbo, GPT-4o: These are developed by OpenAI, excels in dialogue.
Besides, we also compare with several psychological LLMs:
4.2 Evaluation Details
5 Results and Analysis
@@
5.1 Results on Objective Assessments
To explore the performance of different models in our benchmark, we report the results of objective questions in Table 3. There are several observations drawn from the results. First, models with larger parameters generally obtain better results which is consistent with results in many existing studies. For example, YI-1.5-9B demonstrates superior performance compared to YI-1.5-6B. This further illustrates that larger models have a stronger ability in text understanding. Second, when comparing performance between ethics and theoretical knowledge, the results indicate that these models exhibit superior performance in the ethics domain. This discrepancy may be attributed to the closer alignment of ethics aspects with general domain content, suggesting the possibility of the presence of morally relevant data in the training datasets of these general LLMs. Third, within the realm of MCQs, the accuracy for MMCQs is lower than that for SMCQs. This difference is likely due to the increased difficulty in correctly identifying all the correct options in MMCQs. Fourth, our proposed model, fine-tuned on the psychological dataset, demonstrates improvements across most evaluated aspects compared to other models, particularly in terms of standard accuracy, yielding more significant advancements. This underscores the effectiveness of our proposed dataset in enhancing performance within the psychological domain.
During the testing phase, we observed a significant and intriguing pattern in model behavior. GPT-3.5-turbo and GPT-4o frequently provide a single option in MMCQs, whereas Qwen-turbo tends to select all available options.
@@ 5.1 Results on Objective Assessments
5.2 Results on Case-Based QA
We report the results of subjective questions in Table 4. The results obtained from automatic evaluations using ROUGE and BLEU metrics for these models are relatively low. This may be due to the inherent limitations of these metrics in capturing the nuances and complexities of natural language generation, as they primarily focus on surface-level n-gram overlap rather than semantic equivalence and contextual appropriateness. Therefore, we also incorporate BERTScore to evaluate the semantic quality of these models. Most models consistently score between 63 and 65 on the BS indicator, indicating the consistency and stability of the BS evaluation. PsycoLLM achieves the best results among these models on R-1, B-4, and BERTScore, respectively and EmoLLM achieves better results on R-L. Additionally, although EmoLLM and MindChat do not perform well on MCQs, they achieve relatively strong performance in R-1, R-L, and B-4. The main reason might be that these models are fine-tuned with psychological conversation data, enabling them to better grasp case information and generate higher-quality responses.
@@ 5.2 Results on Case-Based QA
5.3 Effect of Difficulty
We further conduct experiments to compare the results of PsycoLLM, Qwen1.5-14B and Qwen-turbo on the Level 2 and Level 3 Psychological Counselor Exams, as detailed in Table 5. Among these models, the performance on theoretical proficiency is generally superior in the Level 3 examination compared to the Level 2 examination across most indicators.
@@ 5.3 Effect of Difficulty
This disparity may be attributed to the Level 2 examination’s requirement for a more extensive understanding of psychological concepts and a deeper comprehension of the relevant theoretical frameworks. However, it is noteworthy that these models exhibit better performance on the professional ethics and case analysis of the Level 2 exam compared to the Level 3 exam. These observation also suggests that under certain circumstances, LLMs may perceive difficulty differently compared to humans. Not all aspects considered complex by humans necessarily pose challenges for models.
6 Related Work
6.1 Benchmark for Mental Health Support
Many work have been proposed for evaluate the performance of the model in various domain. For example, in the general domain, Huang et al. (2024) proposed a Chinese evaluation benchmark to evaluate advanced knowledge and reasoning abilities of Chinese-based models. In the medical domain, Wang et al. (2023) proposed a localized medical benchmark to evaluate the existing LLMs for medicine. In the legal domain, Dai et al. (2023) constructed a Chinese legal benchmark based on the logic of legal practice. In the psychological domain, there are also some studies. Jin et al. (2023) introduced a meticulously crafted benchmark to evaluate the model performance in mental health-related tasks. Zhang et al. (2023c) used GPT-4 to generate question and constructed a psychological benchmark for evaluating the model performance in psychology. Qiu et al. (2023b) created a benchmark with fine-grained labels for each dialogue and focused on dialogue safety in mental health support. However, these benchmark still exist some shortcomings. Some benchmarks narrow their scope to specific subfields or isolated tasks within mental health, neglecting to evaluate the models’ overall psychological knowledge mastery. Furthermore, benchmarks generated by large models carry the risk of deviating from real-world contexts. Compared to these works, we propose a benchmark that incorporates different levels of official examinations in China, enabling a comprehensive assessment of both theoretical knowledge and practical proficiency.
6.2 Mental Application in LLM
The success of LLMs in general domains has sparked interest in their application across various vertical fields. Many LLMs tailored for specific domains have achieved remarkable success, such as HuatuoGPT Zhang et al. (2023b) in the medical domain and DISC-LawLLM Yue et al. (2023) in the legal domain. In the mental health domain, several LLMs also have been proposed to aid help-seekers in accessing effective support from such systems. Sun et al. (2021) proposed a high-quality Chinese dataset of psychological health support in the form of question and answer pair. Qiu et al. (2023a) further used ChatGPT to covert the single-turn to multi-turn dialogue data and subsequently fine-tuned ChatGLM2-6B on this enhanced dataset. Chen et al. (2023b) constructed a multi-turn empathetic conversation dataset and fine-tuned ChatGLM with this data. Xu et al. (2024) fine-tuned the LLM across multiple datasets simultaneously and showed its effectiveness on various mental health prediction tasks. Although these efforts have greatly advanced the field, they rely too heavily on distillations from other models during data generation, leading to the creation of data that may not accurately reflect real-world scenarios. In this paper, we constructed a high-quality psychological dataset, including single-turn, multi-turn dialogue data, and knowledge QA data.
7 Conclusion
@@
Limitations
This study primarily focuses on the development and evaluation of a language model for psychological applications. However, it is important to note that multimodal LLMs, which incorporate both video and speech inputs, are potentially more effective for understanding user emotions due to their ability to process and integrate multiple modalities of data. The creation of multimodal datasets and benchmarks in the psychological domain warrants further investigation. We leave it for future work. Our evaluation was limited to a few selected models as the foundational backbones for fine-tuning. Future research might be interested in expanding this evaluation to include a broader range of models to comprehensively assess their efficacy.
Acknowledgements
The authors would like to express their gratitude to Xubin Li and Xuxiong Liu for their invaluable contributions. Additionally, we extend our thanks to Jingyuan Li, Kaiyuan Xu, Jiaqi Zhao, and Chaoxing Tang for their efforts in model evaluation.
Appendix A Appendix
A.1 Prompts for Multi-turn Dialogue generation
The following prompt is designed to convert single-turn dialogues into a multi-turn dialogue dataset.
This is the prompt for finding evidence.
The following prompt is used for evaluating the quality of the multi-turn data in terms of empathy, supportiveness, guidance, and safety.
The following prompt is used for modifying the multi-turn data according to the above suggestions in terms of empathy, supportiveness, guidance, and safety.

### v2

@@ IV-A Dataset Collection and Preprocessing
In order to accurately evaluate the model performance in the psychology domain, we introduce a benchmark that psychological professional individuals need to master. The data is sourced from publicly available examination questions. We follow a standardized data preprocessing procedure. For some data, we first need to perform optical character recognition (OCR) to convert images to text. After that, we invite several students to manually review the collected data to ensure its quality and consistency with the original document. This process involves rectifying formatting errors, eliminating duplicate questions, and rectifying any instances of garbled characters. Our proposed psychological benchmark draws inspiration from the format of the most authoritative psychological counseling examination in China, comprising two primary components: objective questions and subjective questions. Note that the objective questions mainly comprise two types of multiple-choice questions (MCQs): MCQ with only a single correct option (SMCQ) and MCQ with multiple correct options (MMCQ). The content of benchmark can be categorized into three sections: professional ethics, theoretical proficiency, and case analysis. The professional ethics section generally encompasses general principles such as honesty, integrity, professional responsibility, and social responsibility. Theoretical proficiency typically pertains to fundamental theories and concepts in psychology, including cognitive psychology, developmental psychology, and related fields. Both the professional ethics and theoretical proficiency sections contain only MCQs, designed to evaluate the mastery of knowledge within the psychological domain. For case analysis, it includes both MCQs and iterative questioning based on given case, aiming to better evaluate performance in real-world scenarios. The former focuses on assessing the grasp of professional knowledge related to the case, while the latter offers a more open-ended assessment of the model’s capability in practical utility. Finally, we obtained a total of 3863 MCQs and 20 cases with 100 QAs. An example of MCQs is shown in Fig. 5, and an example of case-based QA is presented in Fig. 6. The statistics of the benchmark are shown in Table III, detailing the distribution of questions across various categories. The questions are classified into three main categories: Level 2 Psychological Counselor Exam, Level 3 Psychological Counselor Exam, and Others. Note that, according to the official definitions, the Level 2 exam is considered more challenging than the Level 3 exam.
IV-B Evaluation Metric
@@
There are three elements in each instance of the case analysis. is the description of the given case, including the background, environment, etc. of the seeker, is the question according to the case, such as the diagnosis, reason and treatment for the case and is the reference answer for each questions. For QA in case analysis, we use the models to generate the response according to the given and . We evaluate the response from model with the corresponding references . We follow existing text generation metric, including ROUGE-1 (R-1), ROUGE-L (R-L) [53], BLEU-4 (B-4) [54], BERTScore [55]. For the ROUGE metric, we report the scores for both R-1 and R-L. R-1 measures unigram overlap, which serves as an indicator of informativeness, while R-L evaluates the longest common subsequence overlap, providing an assessment of fluency. BERTScore leverages embedding from pre-trained encoder to capture the semantic and contextual information of tokens through computing cosine similarities. This metric allows for a more nuanced evaluation of text generation quality by considering the underlying meaning and context rather than just surface-level token matching.
For SMCQ, we use the standard accuracy to evaluate the model performance. For MMCQ, we employ two types of accuracy metrics: standard accuracy and elastic accuracy. Standard accuracy is defined as the ratio of completely correct answers to the total number of questions, providing a clear measure of the model’s precision by counting only fully accurate responses. Elastic accuracy, on the other hand, accounts for partially correct answers that contain no incorrect options but are not entirely complete. This metric offers a evaluation by recognizing the model’s ability to produce answers that are mostly correct, even if they are not fully comprehensive. The elastic accuracy for a MMCQ is computed as follows:
where indicates the count of choices in the answer.
V Experimental Setting
@@
V-A Baseline
In our experiments, our experiment is implemented based on transformers333https://github.com/huggingface/transformers and LLaMA-Factory444https://github.com/hiyouga/LLaMA-Factory.
BaiChuan2555https://github.com/baichuan-inc/Baichuan2: This is a generation of open-source large language models launched by Baichuan Intelligent Technology, which is trained on a high-quality corpus with 2.6 trillion tokens.
ChatGLM3 [24]: This is a generation of pre-trained dialogue models jointly released by Zhipu AI and Tsinghua KEG.
LLaMA3-Chinese [56]666https://github.com/seanzhang-zhichen/llama3-chinese: This model uses LLama3 as the base model and is fine-tuned with high-quality Chinese instruction datasets with the training methods of DORA and LORA+.
Yi (e.g., Yi-1.5-6B, Yi-1.5-9B)777https://github.com/01-ai/Yi: These open-source large language models are trained from scratch by developers at 01.AI, which are trained with a high-quality corpus of 500B tokens.
Qwen (e.g., Qwen1.5-14B-Chat, Qwen-turbo): these are decoder-only transformer models with SwiGLU activation, RoPE, multi-head attention, developed by Alibaba Cloud and are pre-trained using over 3 trillion tokens of high-quality corpus.
GPT-3.5-turbo, GPT-4o: These models, developed by OpenAI and built upon the foundations of GPT-2 and GPT-3, excel in dialogue generation.
Besides, we also compare with several psychological LLMs:
V-B Evaluation Details
@@
VI Results and Analysis
VI-A Results on Objective Assessments
To explore the performance of different models in our benchmark, we report the results of objective questions in Table IV. There are several observations drawn from the results. First, when comparing performance between ethics and theoretical knowledge, the results indicate that these models exhibit superior performance in the ethics domain. This discrepancy may be attributed to the closer alignment of ethics aspects with general domain dataset. It is plausible that the training datasets for these general LLMs include a significant amount of morally relevant data, which enhances their performance in ethical evaluations. In contrast, theoretical knowledge may require more specialized data that is less prevalent in the general-domain training sets. Second, within the realm of MCQs, the accuracy for MMCQs is lower than that for SMCQs. This difference is likely due to the increased difficulty in correctly identifying all the correct options in MMCQs. The requirement to select multiple correct options, as opposed to a single correct option, introduces additional challenges in ensuring comprehensive and accurate responses, thereby affecting overall accuracy. Third, GPT-4o achieves the best results in SMCQ for both theoretical knowledge and case analysis. This demonstrates that this model, despite being trained for general domain, exhibits strong performance in the psychological domain as well. This suggests that its robust general training provides a strong foundation for effectively addressing domain-specific tasks, including those related to psychological concepts and case-based evaluations. Fourth, several models demonstrate the capability to pass the MCQs in the benchmark, primarily extracted from authoritative psychological counseling examinations. Notably, only PsycoLLM achieves an average exceeding 60% in standard accuracy. In terms of elastic accuracy, GPT-4o, Qwen-turbo, Yi-1.5-9B, and PsycoLLM all achieve averages exceeding 60%. This observation highlights the significant potential of using LLM for providing psychological assistance, demonstrating that these models possess the capability to pass relevant exams and deliver reliable support in this domain. Among these models, Yi-1.5-9B is the smallest, yet it still surpasses the 60% in elastic accuracy. FIfth, our proposed model, fine-tuned on the psychological dataset, demonstrates improvements across most evaluated aspects compared to other models, particularly in terms of standard accuracy, yielding more significant advancements. This underscores the effectiveness of our proposed dataset in enhancing performance within the psychological domain.
VI-B Results on Case-Based QA
We report the results of subjective questions in Table V. The results obtained from automatic evaluations using ROUGE and BLEU metrics for these models are relatively low. This may be due to the inherent limitations of these metrics in capturing the nuances and complexities of natural language generation, as they primarily focus on surface-level n-gram overlap rather than semantic equivalence and contextual appropriateness. Therefore, we also incorporate BERTScore into our evaluation framework. Most models consistently score between 63 and 65 on the BS indicator, indicating the consistency and stability of the BS evaluation. PsycoLLM demonstrates superior performance on R-1, B-4, and BERTScore, respectively and EmoLLM achieves better results on R-L. Additionally, although EmoLLM and MindChat do not perform well on MCQs, they achieve relatively strong performance in R-1, R-L, and B-4. The main reason might be that these models are fine-tuned with psychological conversation data, enabling them to better grasp case information and generate higher-quality responses.
We provide a comparative example of the results from PsycoLLM and GPT-4o in Fig. 6. Both models are proficient at identifying the majority of patient issues. A notable distinction emerges in their responses: GPT-4o tends to offer more detailed and specific answers, while PsycoLLM generally provides more concise responses.
VI-C Effect of Difficulty
@@
We further conduct experiments to compare the results of PsycoLLM, Qwen1.5-14B and Qwen-turbo on the Level 2 and Level 3 Psychological Counselor Exams, as detailed in Table VI. Among these models, the performance on theoretical proficiency is generally superior in the Level 3 examination compared to the Level 2 examination across most indicators. This disparity may be attributed to the Level 2 examination’s requirement for a more extensive understanding of psychological concepts and a deeper comprehension of the relevant theoretical frameworks. However, it is noteworthy that these models exhibit better performance on the professional ethics and case analysis of the Level 2 exam compared to the Level 3 exam. These observation also suggests that under certain circumstances, LLMs may perceive difficulty differently compared to humans. Not all aspects considered complex by humans necessarily pose challenges for models.
VI-D Discussion
Model Size
On the one hand, within the same model families, models with larger parameters generally obtain better results in Table IV. For instance, models such as YI-1.5-9B and Qwen1.5-14B-Chat demonstrate better performance compared to their smaller counterparts, YI-1.5-6B and Qwen1.5-7B-Chat, respectively. This performance enhancement can be attributed to the increased capacity of models with larger parameters to capture complex patterns and effectively learn text representations from the data. This observation aligns with findings from a broad spectrum of existing studies that emphasize the advantages of larger model sizes in terms of learning capacity and performance. On the other hand, when compared models from different model families, this findings have slight difference. For example, YI-1.5-9B outperform Baichuan2-13B and Qwen1.5-14B-Chat in the MCQ evaluation. This discrepancy shows that training corpus, and other factors such as training procedures also significantly influence performance.
Model preference in MMCQs
In our experiments, we observe distinct response patterns from different models when addressing MMCQs. For example, GPT-3.5-turbo and GPT-4o frequently provide a single option in MMCQs, whereas Qwen-turbo tends to select all available options. Therefore, GPT-3.5-turbo and GPT-4o have a large improvements when shifting from standard accuracy measure to elastic accuracy metric. The main reason might be that, although GPT-3.5-turbo and GPT-4 are multilingual, they primarily focus on English. In English-speaking contexts, MMCQs are less prevalent compared to SMCQs. Consequently, the training data for these models may predominantly contains SMCQs and lack a substantial MMCQ dataset. This discrepancy leads the models to exhibit a tendency to select only one answer when presented with MMCQs.
Fine-tuned Models vs. General Models
Among these models, EmoLLM, MindChat are fine-tuned with the psychology-related dataset and others are for general domain. Specifically, EmoLLM and MindChat leverage empathetic conversation data for fine-tuning, which contributes to their relatively higher performance in Case-Based QA than general LLMs. However, due to their limited incorporation of knowledge-based data, their performance in MCQ evaluations is average.
VII Conclusion
@@
This study primarily focuses on the development and evaluation of a language model for psychological applications. However, it is important to note that multimodal LLMs, which incorporate both video and speech inputs, are potentially more effective for understanding user emotions due to their ability to process and integrate multiple modalities of data. The creation of multimodal datasets and benchmarks in the psychological domain warrants further investigation. We leave it for future work.